# A Linearly Convergent Frank-Wolfe-type Method for Smooth Convex Minimization over the Spectrahedron

### v1

@@ 3 Convergence Rate Analysis
Moreover, there exists scalars $\mathcal{E}=O\left({\min\left\{{\frac{\alpha{}\delta^{2}}{\beta^{2}},\frac{\delta^{2}}{\beta},\delta,\alpha\lambda_{r^{*}}^{2},\frac{\delta^{3}}{\beta^{2}}}\right\}}\right)$, $\mathcal{E}_{1}=O\left({\frac{\alpha\delta^{2}}{\beta^{2}}}\right)$, $\mathcal{E}_{2}=O(\alpha{}\lambda_{r^{*}}^{2})$ such that on each iteration $t$ in which a drop step is not taken: If $r^{*}=1$ and $h_{t}<\mathcal{E}_{1}$ then
@@
Before proving the theorem let us make a few comments. Result (5) follows in a straightforward manner from the design of the algorithm. Results (6), (7) follow essentially from already known analyses of the Frank-Wolfe method [19] and [14]. The main novel result is the (expected) linear convergence rate in the case $n>r^{*}\geq 2$ given in Eq. (9). Note that the rate in (9) depends explicitly on the rank of optimal solutions $r^{*}$, which is well known to by unavoidable in worst case, see for instance [18]. Also note that the $\frac{\alpha}{\beta}$ term inside the min in the RHS of (9) corresponds (up to a universal constant) to the standard linear convergence rate of (unaccelerated) gradient methods, such as the block Frank-Wolfe method [1]. Finally, Result (8) for the case $r^{*}=n$, i.e., the optimal solutions lie in the relative interior of $\mathcal{S}^{n}$, follows as a simplified case of Result (9).
@@ 3.1 Proof of Theorem 2
The proof of Theorem 2 is constructed from several lemmas that analyze the benefit of each one of the steps employed by Algorithm 1. It is then established how their combination yields the theorem.
The following lemma establishes conditions under-which a drop step will be taken.
@@
Recall that ${\mathbf{X}}_{t+1}^{\textrm{drop}}=\frac{1}{1-c}({\mathbf{X}}_{t}-c{\mathbf{v}}{\mathbf{v}}^{\top})$ for $c=({\mathbf{v}}^{\top}{\mathbf{X}}_{t}^{\dagger}{\mathbf{v}})^{-1}$ and ${\mathbf{v}}\in\textrm{argmax}\{{\mathbf{v}}^{\top}\nabla_{t}{\mathbf{v}}~{}|~{}{\mathbf{v}}\in\textrm{Im}({\mathbf{X}}_{t}),\|{{\mathbf{v}}}\|=1\}$. Using the smoothness of $f$ we have that,
@@
$=f({\mathbf{X}}_{t})+\frac{c}{1-c}\langle{{\mathbf{X}}_{t}-{\mathbf{v}}{\mathbf{v}}^{\top},\nabla_{t}}\rangle+\frac{\beta{}c^{2}}{(1-c)^{2}}.$ (10)
Since $\dim\textrm{Im}({\mathbf{X}}_{t})\geq r^{*}+1$, it follows from the Poincaré separation theorem that ${\mathbf{v}}^{\top}\nabla_{t}{\mathbf{v}}\geq\lambda_{n-r^{*}}(\nabla_{t})=\lambda_{n-r^{*}+1}+\delta_{t}$. On the other-hand, we argue that $\langle{{\mathbf{X}}_{t},\nabla_{t}}\rangle\leq\lambda_{n-r^{*}+1}+\delta_{t}/2$. To see why the latter is true, let us assume by way of contradiction that it is not, and let us denote the matrix ${\mathbf{Y}}_{\eta}=(1-\eta){\mathbf{X}}_{t}+\eta{\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}$ for some $\eta\in[0,1]$ to be specified later, and ${\mathbf{v}}_{t,+}$ as defined in Algorithm 1. Using the smoothness of $f$ again we have that,
$f({\mathbf{Y}}_{\eta})-f^{*}$ $\leq f({\mathbf{X}}_{t})+\eta\langle{{\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}-{\mathbf{X}}_{t},\nabla_{t}}\rangle+\frac{\beta\eta^{2}}{2}\|{{\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}-{\mathbf{X}}_{t}}\|_{F}^{2}$
$\leq f({\mathbf{X}}_{t})-f^{*}-\eta\left({\frac{\delta_{t}}{2}-\eta\beta}\right).$
@@
According to Lemma 2, each drop step reduces the rank of the resulting matrix. On the other-hand, all other steps increase the rank by at most one. Thus, denoting the number of drop steps up to the beginning of some iteration $t$ by $t_{\textrm{drop}}$, it must hold that
@@
Write ${\mathbf{u}}_{t,-}=\tau{\mathbf{z}}+\sqrt{1-\tau^{2}}{\mathbf{w}}$, where ${\mathbf{z}}=\frac{{\mathbf{P}}_{t}^{\perp}{\mathbf{u}}_{t,-}}{\|{{\mathbf{P}}_{t}^{\perp}{\mathbf{u}}_{t,-}}\|}$, $\tau=\|{{\mathbf{P}}_{t}^{\perp}{\mathbf{u}}_{t,-}}\|$, and ${\mathbf{w}}=\frac{{\mathbf{P}}_{t}{\mathbf{u}}_{t,-}}{\|{{\mathbf{P}}_{t}{\mathbf{u}}_{t,-}}\|}$. Consider now some unit vector ${\mathbf{u}}=\tau^{\prime}{\mathbf{z}}+\sqrt{1-\tau^{\prime 2}}{\mathbf{w}}$ with $0\leq\tau^{\prime}\leq\tau$. Fix some step-size $\eta\in[0,1]$ and denote ${\mathbf{Y}}_{\eta}={\mathbf{X}}_{t}+\eta({\mathbf{u}}{\mathbf{u}}^{\top}-{\mathbf{u}}_{t,-}{\mathbf{u}}_{t,-}^{\top})$. Using the smoothness of $f$ it holds that,
@@
$\leq f({\mathbf{X}}_{t})-\eta(\tau^{2}-\tau^{\prime 2})\delta_{t}+\eta^{2}\beta\left({1-(\tau\tau^{\prime}+\sqrt{1-\tau^{2}}\sqrt{1-\tau^{\prime 2}})^{2}}\right)$
@@
$\underset{(a)}{\leq}f({\mathbf{X}}_{t})-\eta(\tau^{2}-\tau^{\prime 2})\delta_{t}+\eta^{2}\beta\left({\tau^{2}+\tau^{\prime 2}-2\tau^{2}\tau^{\prime 2}-2\tau\tau^{\prime}(1-\tau^{2})}\right)$
@@
where (a) follows from the assumption $\tau^{\prime}\leq\tau$.
@@
Using the feasible step-size $\eta^{*}=\min\{\frac{\delta_{t}}{2\beta},\gamma_{t}\}$ then gives,
@@
Consider now the case $\tau^{\prime}=(1-c)\tau$ for some scalar $c\in[0,1]$ to be specified in the sequel. Using the smoothness of $f$ again, the definition of ${\mathbf{u}}_{t,+}$ in Algorithm 1 (see also Eq. (4)), and applying (3.1) with $\eta=\gamma_{t}$ we have that,
@@
where the last inequality follows from plugging-in (13).
@@
$\leq\textrm{Tr}({\mathbf{P}}_{t}^{\perp}{\mathbf{U}}_{t}{\mathbf{U}}_{t}^{\top})=\textrm{Tr}\left({{\mathbf{P}}_{t}^{\perp}\mathbb{E}_{t}[r_{t}{\mathbf{u}}_{t,-}{\mathbf{u}}_{t,-}^{\top})]}\right)=\mathbb{E}_{t}[r_{t}\textrm{Tr}\left({{\mathbf{P}}_{t}^{\perp}{\mathbf{u}}_{t,-}{\mathbf{u}}_{t,-}^{\top})}\right]$
$=r_{t}\mathbb{E}_{t}[\|{{\mathbf{P}}_{t}^{\perp}{\mathbf{u}}_{t,-}}\|^{2}].$
The following lemma is central to the analysis of the benefit of both standard Frank-Wolfe steps and away steps in Algorithm 1. It lower-bounds the eigengap between the extreme eigenvalues associated with the principal $r^{*}$-dimensional subspace (corresponding to the $r^{*}$ lowest eigenvalues) of the current gradient direction $\nabla_{t}$. This is essentially a type of a Polyak-Łojasiewicz (PL) inequality w.r.t. to the face of $\mathcal{S}^{n}$ corresponding to the this principal subspace of $\nabla_{t}$.
@@
It holds that,
@@
=$\underset{(c)}{=}\sqrt{r^{*}}\|{{\mathbf{X}}_{t}-{\mathbf{X}}^{*}}\|_{F}\left({\lambda_{n-r^{*}+1}(\nabla_{t})-\lambda_{n}(\nabla_{t})}\right)-\textrm{Tr}({\mathbf{W}}^{\perp})\lambda_{n}(\nabla_{t})$
@@
It also holds that,
$\langle{{\mathbf{W}}^{\perp},\nabla_{t}}\rangle$ $=\langle{{\mathbf{P}}_{t}^{\perp}({\mathbf{X}}_{t}-{\mathbf{X}}^{*}){\mathbf{P}}_{t}^{\perp},\nabla_{t}}\rangle$
$\leq\textrm{Tr}({\mathbf{P}}_{t}^{\perp}{\mathbf{X}}_{t})\lambda_{1}(\nabla_{t})-\textrm{Tr}({\mathbf{P}}_{t}^{\perp}{\mathbf{X}}^{*})\lambda_{n-r^{*}}(\nabla_{t}).$ (16)
@@
Rearraning the inequality, using the assumption on $\textrm{Tr}({\mathbf{P}}_{t}^{\perp}{\mathbf{X}}_{t})$, and the quadratic growth bound $\|{{\mathbf{X}}_{t}-{\mathbf{X}}^{*}}\|_{F}\leq\sqrt{\frac{2h_{t}}{\alpha}}$, yields the lemma.
The following Lemma, which builds on the result of the previous Lemma 6, establishes the reduction in objective function value for FW / away steps, under a positive in-face eigengap.
Lemma 7 (FW/away steps under in-face eignegap).
@@
Since $\textrm{rank}({\mathbf{X}}_{t})\geq r^{*}$, according to the Poincaré separation theorem, we have that there exists a unit vector ${\mathbf{v}}\in\textrm{Im}({\mathbf{X}}_{t})$ such that ${\mathbf{v}}^{\top}\nabla_{t}{\mathbf{v}}\geq\lambda_{n-r^{*}+1}(\nabla_{t})$, which in turn implies that ${\mathbf{v}}_{t,-}^{\top}\nabla_{t}{\mathbf{v}}_{t,-}\geq\lambda_{n-r^{*}+1}(\nabla_{t})$. Also, we clearly have that ${\mathbf{v}}_{t,+}^{\top}\nabla_{t}{\mathbf{v}}_{t,+}=\lambda_{n}(\nabla_{t})$. This gives,
@@
where in the last inequality we have used the fact that $\eta\in[0,1)$ and $\langle{{\mathbf{X}}_{t}-{\mathbf{v}}_{t,-}{\mathbf{v}}_{t,-}^{\top},\nabla_{t}}\rangle\geq 0$.
@@
which proves the lemma.

### v2

@@ 3 Convergence Rate Analysis
Moreover, there exist scalars $\mathcal{E}=O\left({\min\left\{{\frac{\delta^{2}}{\beta},\delta,\alpha\lambda_{r^{*}}^{2},\frac{\delta^{3}}{\beta^{2}}}\right\}}\right)$, $\mathcal{E}_{1}=O\left({\frac{\delta^{2}}{\beta}}\right)$, $\mathcal{E}_{2}=O(\alpha\lambda_{r^{*}}^{2})$ such that on each iteration $t$ in which a drop step is not taken:
If $r^{*}=1$ and $h_{t}<\mathcal{E}_{1}$ then
@@
Before proving the theorem let us make a few comments. Result (6) follows in a straightforward manner from the design of the algorithm. Results (7), (8) follow essentially from already known analyses of the Frank-Wolfe method [21] and [15]. The main novel result is the (expected) linear convergence rate in the case $n>r^{*}\geq 2$ given in Eq. (10). Note that the rate in (10) depends explicitly on the rank of optimal solutions $r^{*}$, which is well-known to be unavoidable in worst case, see for instance [20]. Also note that the $\frac{\alpha}{\beta}$ term inside the min in the RHS of (10) corresponds (up to a universal constant) to the standard linear convergence rate of (unaccelerated) gradient methods, such as the block Frank-Wolfe method [1]. Finally, Result (9) for the case $r^{*}=n$, i.e., the optimal solutions lie in the relative interior of $\mathcal{S}^{n}$, follows as a simplified case of Result (10).
@@ 3.1 Proof of Theorem 2
Roadmap for the proof:
As mentioned above, the most challenging part is the proof of Result (10). We consider the run of Algorithm 1 in two phases. The first one is a finite burn-in phase, before the strict complementarity condition has been “detected”, and the standard FW steps are used to drive a provable sublinear convergence rate (Result (7)). Once the approximation error $h_{t}$ reaches a critical value (which naturally scales with the strict complementarity measure $\delta$), we enter the second phase of the run in which the convergence is linear, in expectation. In this second phase, the algorithm automatically adapts to the optimal rank $r^{*}$ — on each iteration $t$ in which a drop step is not taken, it must hold that $\textrm{rank}({\mathbf{X}}_{t})\leq r^{*}$ (Lemma 3). On each such iteration we then consider two cases. In the first case, the current iterate ${\mathbf{X}}_{t}$ is not sufficiently aligned with the principal subspace of the gradient $\nabla{}f({\mathbf{X}}_{t})$, in the sense that $\textrm{Tr}({\mathbf{P}}_{t}^{\perp}{\mathbf{X}}_{t})>c{}h_{t}$ for some constant $c$. In this case, Lemmas 4 and 5 guarantee that a pairwise step will reduce the approximation error, in expectation, by $\Omega(h_{t})$ (compatible with a linear convergence rate, in expectation). Here it is important that $\textrm{rank}({\mathbf{X}}_{t})\leq r^{*}$ which, together with the quadratic growth property, guarantees the pairwise step can be taken with sufficiently large step-size. In the other case, the iterate ${\mathbf{X}}_{t}$ is sufficiently aligned with the principal subspace of $\nabla{}f({\mathbf{X}}_{t})$, and we can use the quadratic growth of $f$ to establish a Polyak-Łojasiewicz-type (PL) inequality (a well-known property used to derive linear convergence rates for first-order methods) on the face of $\mathcal{S}^{n}$ corresponding to this subspace (Lemma 6). With this inequality in place, we can establish that either a standard FW step or an away step will lead to a $\Omega(h_{t})$ (deterministic) reduction in approximation error, which is again compatible with a linear convergence rate.
The following lemma establishes conditions under which a drop step will be taken.
@@
Recall that ${\mathbf{X}}_{t+1}^{\textrm{drop}}=\frac{1}{1-c}({\mathbf{X}}_{t}-c{\mathbf{v}}{\mathbf{v}}^{\top})$ for $c=({\mathbf{v}}^{\top}{\mathbf{X}}_{t}^{\dagger}{\mathbf{v}})^{-1}$ and that ${\mathbf{v}}\in\textrm{argmax}\{{\mathbf{v}}^{\top}\nabla_{t}{\mathbf{v}}~|~{\mathbf{v}}\in\textrm{Im}({\mathbf{X}}_{t}),\|{{\mathbf{v}}}\|=1\}$. Using the smoothness of $f$ we have that,
@@
$\leq f({\mathbf{X}}_{t})+\frac{c}{1-c}\langle{{\mathbf{X}}_{t}-{\mathbf{v}}{\mathbf{v}}^{\top},\nabla_{t}}\rangle+\frac{\beta{}c^{2}}{(1-c)^{2}}.$ (11)
Since $\dim\textrm{Im}({\mathbf{X}}_{t})\geq r^{*}+1$, it follows from the Poincaré inequality (e.g., Theorem 4.4 in [3]) that ${\mathbf{v}}^{\top}\nabla_{t}{\mathbf{v}}\geq\lambda_{n-r^{*}}(\nabla_{t})=\lambda_{n-r^{*}+1}(\nabla_{t})+\delta_{t}$ [^6] Poincaré inequality states that for any ${\mathbf{A}}\in\mathbb{S}^{n}$ and a subspace $\mathcal{V}\subseteq\mathbb{R}^{n}$ of dimension $k$, there must exist a unit vector ${\mathbf{w}}\in\mathcal{V}$ such that ${\mathbf{w}}^{\top}{\mathbf{A}}{\mathbf{w}}\geq\lambda_{n-k+1}({\mathbf{A}})$.. On the other hand, we argue that $\langle{{\mathbf{X}}_{t},\nabla_{t}}\rangle\leq\lambda_{n-r^{*}+1}(\nabla_{t})+\delta_{t}/2$. To see why the latter is true, let us assume by way of contradiction that it is not, and let us denote the matrix ${\mathbf{Y}}_{\eta}=(1-\eta){\mathbf{X}}_{t}+\eta{\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}$ for some $\eta\in[0,1]$ to be specified later, and ${\mathbf{v}}_{t,+}$ as defined in Algorithm 1. Using the smoothness of $f$ again we have that,
$f({\mathbf{Y}}_{\eta})-f^{*}$ $\leq f({\mathbf{X}}_{t})+\eta\langle{{\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}-{\mathbf{X}}_{t},\nabla_{t}}\rangle+\frac{\beta\eta^{2}}{2}\|{{\mathbf{v}}_{t,+}{\mathbf{v}}_{t,+}^{\top}-{\mathbf{X}}_{t}}\|_{F}^{2}-f^{*}$
$\leq f({\mathbf{X}}_{t})-f^{*}+\eta\left({\lambda_{n}(\nabla_{t})-\lambda_{n-r^{*}+1}(\nabla_{t})-\delta_{t}/2}\right)+\eta^{2}\beta$
$\leq f({\mathbf{X}}_{t})-f^{*}-\eta\left({\frac{\delta_{t}}{2}-\eta\beta}\right),$
where the middle inequality is due to the definition of ${\mathbf{v}}_{t,+}$ in the algorithm and our assumption by way of contradiction that $\langle{{\mathbf{X}}_{t},\nabla_{t}}\rangle>\lambda_{n-r^{*}+1}(\nabla_{t})+\delta_{t}/2$.
@@
According to Lemma 2, each drop step reduces the rank of the resulting matrix. On the other hand, all other steps increase the rank by at most one. Thus, denoting the number of drop steps up to the beginning of some iteration $t$ by $t_{\textrm{drop}}$, it must hold that
@@
Write ${\mathbf{u}}_{t,-}$ as ${\mathbf{u}}_{t,-}=\tau{\mathbf{z}}+\sqrt{1-\tau^{2}}{\mathbf{w}}$, where ${\mathbf{z}}=\frac{{\mathbf{P}}_{t}^{\perp}{\mathbf{u}}_{t,-}}{\|{{\mathbf{P}}_{t}^{\perp}{\mathbf{u}}_{t,-}}\|}$, $\tau=\|{{\mathbf{P}}_{t}^{\perp}{\mathbf{u}}_{t,-}}\|$, and ${\mathbf{w}}=\frac{{\mathbf{P}}_{t}{\mathbf{u}}_{t,-}}{\|{{\mathbf{P}}_{t}{\mathbf{u}}_{t,-}}\|}$. Note that by definition of ${\mathbf{P}}_{t}$ and ${\mathbf{P}}_{t}^{\perp}$ we have that
${\mathbf{z}}^{\top}\nabla_{t}{\mathbf{z}}-{\mathbf{w}}^{\top}\nabla_{t}{\mathbf{w}}\geq\lambda_{n-r^{*}}(\nabla_{t})-\lambda_{n-r^{*}+1}(\nabla_{t})=\delta_{t}.$ (13)
Consider now some unit vector ${\mathbf{u}}=\tau^{\prime}{\mathbf{z}}+\sqrt{1-\tau^{\prime 2}}{\mathbf{w}}$ with $0\leq\tau^{\prime}\leq\tau$. Fix some step-size $\eta\in[0,1]$ and denote ${\mathbf{Y}}_{\eta}={\mathbf{X}}_{t}+\eta({\mathbf{u}}{\mathbf{u}}^{\top}-{\mathbf{u}}_{t,-}{\mathbf{u}}_{t,-}^{\top})$. Using the smoothness of $f$ it holds that,
@@
$\underset{(a)}{\leq}f({\mathbf{X}}_{t})-\eta(\tau^{2}-\tau^{\prime 2})\delta_{t}+\eta^{2}\beta\left({1-(\tau\tau^{\prime}+\sqrt{1-\tau^{2}}\sqrt{1-\tau^{\prime 2}})^{2}}\right)$
@@
$\underset{(b)}{\leq}f({\mathbf{X}}_{t})-\eta(\tau^{2}-\tau^{\prime 2})\delta_{t}+\eta^{2}\beta\left({\tau^{2}+\tau^{\prime 2}-2\tau^{2}\tau^{\prime 2}-2\tau\tau^{\prime}(1-\tau^{2})}\right)$
@@
where (a) is due to Eq. (13) and recalling that $\tau^{\prime}\in[0,\tau]$, and (b) also follows since $\tau^{\prime}\leq\tau$.
@@
Using the step-size $\eta^{*}=\min\{\frac{\delta_{t}}{2\beta},\gamma_{t}\}$, which satisfies that ${\mathbf{Y}}_{\eta^{*}}\in\mathcal{S}^{n}$ (see Lemma 2), then gives
@@
Consider now the case $\tau^{\prime}=(1-c)\tau$ for some scalar $c\in[0,1]$ to be specified in the sequel. Using the smoothness of $f$ again, the definition of ${\mathbf{u}}_{t,+}$ in Algorithm 1 (see also Eq. (4)), and applying (3.1) with $\eta=\gamma_{t}$ (note that by definition $\gamma_{t}\in(0,1]$) we have that,
@@
where the last inequality follows from plugging (15) into the term $2\gamma_{t}\beta\tau^{2}$.
@@
$\leq\textrm{Tr}({\mathbf{P}}_{t}^{\perp}{\mathbf{U}}_{t}{\mathbf{U}}_{t}^{\top})\underset{(a)}{=}\textrm{Tr}\left({{\mathbf{P}}_{t}^{\perp}\mathbb{E}_{t}[r_{t}{\mathbf{u}}_{t,-}{\mathbf{u}}_{t,-}^{\top}]}\right)=\mathbb{E}_{t}\left[{r_{t}\textrm{Tr}\left({{\mathbf{P}}_{t}^{\perp}{\mathbf{u}}_{t,-}{\mathbf{u}}_{t,-}^{\top}}\right)}\right]$
$=r_{t}\mathbb{E}_{t}[\|{{\mathbf{P}}_{t}^{\perp}{\mathbf{u}}_{t,-}}\|^{2}],$
where (a) holds since by definition of ${\mathbf{u}}_{t,-}$ in Algorithm 1 we have that $\mathbb{E}_{t}[{\mathbf{u}}_{t,-}{\mathbf{u}}_{t,-}^{\top}]=\frac{1}{r_{t}}{\mathbf{U}}_{t}{\mathbf{U}}_{t}^{\top}$. ∎
The following lemma is central to the analysis of the benefit of both standard Frank-Wolfe steps and away steps in Algorithm 1. It lower-bounds the eigengap between the extreme eigenvalues associated with the principal $r^{*}$-dimensional subspace (corresponding to the $r^{*}$ lowest eigenvalues) of the current gradient direction $\nabla_{t}$. This is essentially a type of Polyak-Łojasiewicz (PL) inequality w.r.t. the face of $\mathcal{S}^{n}$ corresponding to the principal subspace of $\nabla_{t}$.
@@
Note that by definition of ${\mathbf{P}}_{t}$ and $\mathbf{W}$}, for any eigenvector $\mathbf{u}$} of $\mathbf{W}$} corresponding to a non-zero eigenvalue, we have that
$\lambda_{n}(\nabla_{t})=\min_{{\mathbf{v}}:{\mathbf{P}}_{t}{\mathbf{v}}={\mathbf{v}},\|{{\mathbf{v}}}\|=1}{\mathbf{v}}^{\top}\nabla_{t}{\mathbf{v}}\leq{\mathbf{u}}^{\top}\nabla_{t}{\mathbf{u}}\leq\max_{{\mathbf{v}}:{\mathbf{P}}_{t}{\mathbf{v}}={\mathbf{v}},\|{{\mathbf{v}}}\|=1}{\mathbf{v}}^{\top}\nabla_{t}{\mathbf{v}}=\lambda_{n-r^{*}+1}(\nabla_{t}).$ (17)
Thus, we have that,
@@
$\underset{(c)}{=}\sqrt{r^{*}}\|{{\mathbf{X}}_{t}-{\mathbf{X}}^{*}}\|_{F}\left({\lambda_{n-r^{*}+1}(\nabla_{t})-\lambda_{n}(\nabla_{t})}\right)-\textrm{Tr}({\mathbf{W}}^{\perp})\lambda_{n}(\nabla_{t})$
@@
Similarly to Eq. (17), by the definition of ${\mathbf{P}}_{t}^{\perp}$, we have that
$\lambda_{n-r^{*}}(\nabla_{t})=\min_{{\mathbf{v}}:{\mathbf{P}}_{t}^{\perp}{\mathbf{v}}={\mathbf{v}},\|{{\mathbf{v}}}\|=1}{\mathbf{v}}^{\top}\nabla_{t}{\mathbf{v}}\leq\max_{{\mathbf{v}}:{\mathbf{P}}_{t}^{\perp}{\mathbf{v}}={\mathbf{v}},\|{{\mathbf{v}}}\|=1}{\mathbf{v}}^{\top}\nabla_{t}{\mathbf{v}}=\lambda_{1}(\nabla_{t}).$
Thus, using the definition of $\mathbf{W}$$^{\perp}$, we have that
$\langle{{\mathbf{W}}^{\perp},\nabla_{t}}\rangle$ $=\langle{{\mathbf{P}}_{t}^{\perp}{\mathbf{X}}_{t}{\mathbf{P}}_{t}^{\perp},\nabla_{t}}\rangle-\langle{{\mathbf{P}}_{t}^{\perp}{\mathbf{X}}^{*}{\mathbf{P}}_{t}^{\perp},\nabla_{t}}\rangle$
$\leq\textrm{Tr}({\mathbf{P}}_{t}^{\perp}{\mathbf{X}}_{t})\lambda_{1}(\nabla_{t})-\textrm{Tr}({\mathbf{P}}_{t}^{\perp}{\mathbf{X}}^{*})\lambda_{n-r^{*}}(\nabla_{t}),$ (19)
where the inequality uses the fact that ${\mathbf{P}}_{t}^{\perp}{\mathbf{X}}_{t}{\mathbf{P}}_{t}^{\perp}$ and ${\mathbf{P}}_{t}^{\perp}{\mathbf{X}}^{*}{\mathbf{P}}_{t}^{\perp}$ are positive semidefinite.
@@
Rearranging the inequality, using the assumption on $\textrm{Tr}({\mathbf{P}}_{t}^{\perp}{\mathbf{X}}_{t})$, and the quadratic growth bound $\|{{\mathbf{X}}_{t}-{\mathbf{X}}^{*}}\|_{F}\leq\sqrt{\frac{2h_{t}}{\alpha}}$, yields the lemma.
The following Lemma, which is motivated by the result of the previous Lemma 6, establishes the reduction in objective function value for FW / away steps, under a positive in-face eigengap.
Lemma 7 (FW/away steps under in-face eigengap).
@@
Since $\textrm{rank}({\mathbf{X}}_{t})\geq r^{*}$, according to the Poincaré inequality (e.g., Theorem 4.4 in [3], see also Footnote 6), we have that there exists a unit vector ${\mathbf{v}}\in\textrm{Im}({\mathbf{X}}_{t})$ such that ${\mathbf{v}}^{\top}\nabla_{t}{\mathbf{v}}\geq\lambda_{n-r^{*}+1}(\nabla_{t})$, which in turn implies that ${\mathbf{v}}_{t,-}^{\top}\nabla_{t}{\mathbf{v}}_{t,-}\geq\lambda_{n-r^{*}+1}(\nabla_{t})$. Also, we clearly have that ${\mathbf{v}}_{t,+}^{\top}\nabla_{t}{\mathbf{v}}_{t,+}=\lambda_{n}(\nabla_{t})$. This gives,
@@
where in the last inequality we have used the fact that $\eta\in[0,1)$ and $\langle{{\mathbf{X}}_{t}-{\mathbf{v}}_{t,-}{\mathbf{v}}_{t,-}^{\top},\nabla_{t}}\rangle\leq 0$.
@@
where we have used the fact that with our choice of $\eta$, which satisfies $\eta\leq 1/2$, we have that $\frac{\eta^{2}\beta}{(1-\eta)^{2}}\leq\frac{\eta^{2}\beta}{(1-1/2)^{2}}=\frac{\delta_{t,r^{*}}^{2}}{64\beta}$. ∎